a sample of cobalt-60 whose original activity is 64 Bq, how long would it take to decay to activity of 1Bq
1
Expert's answer
2021-02-25T11:26:23-0500

Let’s use the formula for radioactive decay:


"A=A_0e^{-\lambda t},""\dfrac{A}{A_0}=e^{-\lambda t},""ln(\dfrac{A}{A_0})=ln(e^{-\lambda t}),""ln(\dfrac{A}{A_0})=-\dfrac{0.693}{T_{1/2}}t,""t=-\dfrac{ln(\dfrac{A}{A_0})}{0.693}T_{1/2}=-\dfrac{ln(\dfrac{1\ Bq}{64\ Bq})}{0.693}\cdot5.27\ yr=31.6\ yr."